OpenStudy (anonymous):

Graph y=3cos(2x-(pi/2))

OpenStudy (calculusxy):

How do u graph y=3cos(2(x+pi/4))+2?? ** y=3cos(2(x+pi/4))+2 Equation for graphing sin function: y=Acos(Bx-C), with A=amplitude, Period=2π/B, Phase shift=C/B. For given sin function: y=3cos(2(x+pi/4))+2=3cos(2x+pi/2))+2 Amplitude=3 B=2 Period=2π/B=2π/2=π 1/4 period=π/4 Phase shift=C/B=(π/2)/2=π/4 (shift to the left) Curve is shifted vertically 2 units y-intercept set x=0 y=3cos(2x+pi/2))+2=3cos(π/2)+2=0+2=2 .. Graphing for one period: Without any phase shift or vertical shift, you can plot the given cos curve (with amplitude=3) on an (x,y) coordinate system with the following coordinates: (0,3), (π/4,0), (π/2,-3), (3π/4,0), (π,3) With a phase shift of π/4 to the left, the x-coordinates change to read as follows: (-π/4,3), (0,0), (π/4,-3), (π/2,0), (3π/4,3) Adding a vertical shift of 2 units cause the y-coordinates to change, resulting in the final configuration of the graph as follows: (-π/4,5), (0,2), (π/4,-1), (π/2,2), (3π/4,5) Source: algebra.com

OpenStudy (anonymous):

A plot is attached.

OpenStudy (anonymous):

Ah I see, I know what I have to do now. Thanks!